# GARCH copulas and GARCH-mimicking copulas

### v1

@@ 4.2 V-transformation and inverse v-transformation of copulas
The following result, which is an adaptation of a result in McNeil, (2021), is key to constructing new copulas using v-transforms and their inverses.
@@
Let $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ be v-transforms with associated partial inverses $\mathcal{V}_{1}^{-1}$ and $\mathcal{V}_{2}^{-1}$ and $\Delta$-functions $\Delta_{1}$ and $\Delta_{2}$. Let $(U,V)^{\top}$ be distributed according to a copula $C_{U,V}$ with density $c_{U,V}$.
@@
    $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$ is distributed according to a copula $C_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ with density
    $c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)=\;$ $c_{U,V}(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v))\;\Delta_{1}(u)\,\Delta_{2}(v)\;+$
    $c_{U,V}(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v))\;\big{(}1-\Delta_{1}(u)\big{)}\,\Delta_{2}(v)\;+$ (14)
    $c_{U,V}(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v)+v)\;\Delta_{1}(u)\,\big{(}1-\Delta_{2}(v)\big{)}\;+$
    $c_{U,V}(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v)+v)\;\big{(}1-\Delta_{1}(u)\big{)}\,\big{(}1-\Delta_{2}(v)\big{)}.$
@@
    $(\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V))^{\top}$ is distributed according to a copula $C_{\mathcal{V}^{\leftarrow}_{1}(U),\mathcal{V}^{\leftarrow}_{2}(V)}$ with density
    $c_{\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V)}(u,v)=c_{U,V}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v)).$ (15)
@@
We refer to the copula density $c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ in part (i) as a v-transformed version of $c_{U,V}$ while we refer to the copula density $c_{\mathcal{V}_{1}^{\leftarrow}(U),\mathcal{V}_{2}^{\leftarrow}(V)}$ in part (ii) as an inverse-v-transformed version of $c_{U,V}$. Inverse v-transformation of a copula density using (15) followed by v-transformation using (14) returns the original copula density $c_{U,V}$ (since for any v-transform $\mathcal{V}\circ\mathcal{V}^{\leftarrow}(y)=y$ for $y\in[0,1]$). However v-transformation of a copula density using (14) followed by inverse v-transformation using (15) does not, in general, return the original density $c$, except in special cases, such as the following.
@@
If $C_{U,V}$ is a jointly symmetric copula and $\mathcal{V}_{1}(u)=\mathcal{V}_{2}(u)=|2u-1|$ for $u\in[0,1]$, then the density of the copula $C_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}$ satisfies
$c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)$ $=$ $c_{U,V}\left(\frac{1-u}{2},\frac{1-v}{2}\right)=c_{U,V}\left(\frac{1-u}{2},\frac{1+v}{2}\right)$ (16)
$=$ $c_{U,V}\left(\frac{1+u}{2},\frac{1-v}{2}\right)=c_{U,V}\left(\frac{1+u}{2},\frac{1+v}{2}\right)$
and
$c_{\mathcal{V}_{1}^{\leftarrow}\circ\mathcal{V}_{1}(U),\mathcal{V}_{2}^{\leftarrow}\circ\mathcal{V}_{2}(V)}(u,v)=c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(|2u-1|,|2v-1|)=c_{U,V}(u,v).$ (17)
@@
By part (i) of Theorem 1 and equation (14), the copula density satisfies
$c_{\mathcal{V}_{1}(U),\mathcal{V}_{2}(V)}(u,v)$ $=\frac{1}{4}\left(c_{U,V}\left(\frac{1-u}{2},\frac{1-v}{2}\right)+c_{U,V}\left(\frac{1-u}{2},\frac{1+v}{2}\right)\right.$
$\left.+c_{U,V}\left(\frac{1+u}{2},\frac{1-v}{2}\right)+c_{U,V}\left(\frac{1+u}{2},\frac{1+v}{2}\right)\right)$
and (16) follows easily from the joint symmetry property of $C_{U,V}$. The identity (17) is then easily verified by considering the values of $|2u-1|$ and $|2v-1|$ in the four quadrants of the unit square defined by the lines $u=0.5$ and $v=0.5$ and using each of the four equations in (16) in turn. ∎
Remark 2.
We can easily derive a corresponding result for the simpler case where the copula $C_{U,V}$ has h-symmetry and $V$ is v-transformed with the symmetric v-transform. The copula density of the random vector $(U,|2V-1|))^{\top}$ can be shown to satisfy
$c_{U,|2V-1|}(u,v)=c_{U,V}\left(u,\frac{1-v}{2}\right)=c_{U,V}\left(u,\frac{1+v}{2}\right)$ (18)
and it is easily verified that $c_{U,|2V-1|}(u,|2v-1|)=c_{U,V}(u,v)$.
Remark 3.
To see that the identity (17) does not always hold when $\mathcal{V}_{1}(u)=\mathcal{V}_{2}(u)=|2u-1|$, it suffices to start with any copula $C_{U,V}$ that is not jointly symmetric, since application of the inverse v-transform in the final step always results in a jointly symmetric copula. For example, when $C_{U,V}=C^{\text{Ga}}_{\rho}$, the Gaussian copula with parameter $\rho>0$, v-transformation followed by inverse v-transformation results in the jointly symmetric mixture $0.5C^{\text{Ga}}_{\rho}+0.5C^{\text{Ga}}_{-\rho}$.
Corollary 1 and Remark 2 allow us to state the following simple result, which is relevant to our study of the symmetries in the densities of copulas of GARCH(1,1) type.
Proposition 6.
Let $C$ be an absolutely continuous copula with density $c$.
  1. (i)
    $C$ is h-symmetric if and only if $c(u,v)=c^{*}(u,|2v-1|)$ for some copula density $c^{*}$.
  2. (ii)
    $C$ is jointly symmetric if and only if $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ for some copula density $c^{*}$.
Proof.
It is trivial to verify that copulas with the structures $c^{*}(u,|2v-1|)$ and $c^{*}(|2u-1|,|2v-1|)$ are h-symmetric and jointly symmetric, respectively. To show the reverse implication in part (i), we choose $c^{*}$ to be the copula density of $(U,|2V-1|)^{\top}$ where $(U,V)^{\top}$ is a random vector with copula density $c$. To show the reverse implication in part (ii), we choose $c^{*}$ to be the copula density of $(|2U-1|,|2V-1|)^{\top}$ where $(U,V)^{\top}$ is a random vector with copula density $c$. ∎
@@ 4.3 Implicit v-transforms in general bivariate densities
In this section we use part (i) of Theorem 1 to find a condition for our generic random vector $(Y,Z)^{\top}$ on $\mathbb{R}^{2}$ which allows us to write its copula density $c$ in the general form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ for v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ and a copula density $c^{*}$. This result requires the concept of a v-shaped transformation on $\mathbb{R}$, which should be distinguished from a v-transform.
@@
A function $T$ is a v-shaped function on $\mathbb{R}$ if it can be written in the form
$T(x)=\begin{cases}T_{L}(\mu_{T}-x)&\quad x\leqslant\mu_{T}\\ T_{R}(x-\mu_{T})&\quad x>\mu_{T}\end{cases}$
where $T_{L}$ and $T_{R}$ are strictly increasing, continuous and differentiable functions on $\mathbb{R}^{+}=[0,\infty)$ such that $T_{L}(0)=T_{R}(0)$ and $\mu_{T}\in\mathbb{R}$.
@@
$\frac{f_{Y,Z}(y,z)}{f_{Y}(y)f_{Z}(z)}=g\left(T_{1}(y),T_{2}(z)\right)$
for two v-shaped transformations $T_{1}$ and $T_{2}$ on $\mathbb{R}$ and some function $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}^{+}$, then we may write $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ on $[0,1]^{2}$ where $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are v-transforms given by $\mathcal{V}_{1}(u)=F_{T_{1}(Y)}\circ T_{1}\circ F_{Y}^{-1}(u)$ and $\mathcal{V}_{2}(v)=F_{T_{2}(Z)}\circ T_{2}\circ F_{Z}^{-1}(v)$ and where $c^{*}$ is the density of the copula of $(T_{1}(Y),T_{2}(Z))^{\top}$.
@@
We give two examples to illustrate the use of Theorem 2.
Example 1 (Symmetric GARCH-type process).
Theorem 2 gives an alternative proof of the fact that the copula density $c_{1}$ in a symmetric GARCH-type process can be written in the form $c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of the copula of $(|X_{1}|,|X_{2}|)^{\top}$. The joint density in (6) may be written as
$f_{X_{1},X_{2}}(x,y)=\int_{0}^{\infty}\frac{1}{w\sigma\left(|x|,s\right)}f_{\epsilon}\left(\frac{|y|}{\sigma\left(|x|,s\right)}\right)f_{\epsilon}\left(\frac{|x|}{s}\right)f_{\sigma}(s)\mathrm{d}s$
in terms of the absolute value function, which is a v-shaped transformation on $\mathbb{R}$; moreover the marginal distribution satisfies $f_{X}(x)=f_{X}(|x|)$.
@@
and conclude that the copula density satisfies $c_{1}(u,v)=c_{1}^{*}(\mathcal{V}_{1}(u),|2v-1|)$ where $c_{1}^{*}$ is the copula density of $(\sigma(X_{1}),|X_{2}|)^{\top}$ or, in other words, $(\sigma_{t},|X_{t}|)^{\top}$ for any $t$. To derive $\mathcal{V}_{1}$ we must find the composition of functions $\mathcal{V}_{1}(u)=F_{\sigma}\circ\sigma\circ F_{X}^{-1}(u)$. By using the fact that the distribution function of $\sigma_{t}$ is
$F_{\sigma}(s)=F_{X}\left(\sqrt{\frac{s^{2}-\alpha_{0}}{\alpha_{1}}}\right)-F_{X}\left(-\sqrt{\frac{s^{2}-\alpha_{0}}{\alpha_{1}+\gamma_{1}}}\right),$
we obtain after some calculation that
@@
For the classical ARCH(1) process with asymmetric innovation density in the upper right panel of Figure 3 it does not seem to be generally possible to write the copula density in the form $c_{1}(u,v)=c_{1}^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$. Although the density ratio can be written in the form
$\frac{f_{X_{1},X_{2}}(x,y)}{f_{X}(x)f_{X}(y)}=\frac{f_{\epsilon}\left(\frac{y}{\sigma(x)}\right)}{\sigma(x)f_{X}(y)}=h(\sigma(x),y),$
it does not appear to be possible to find a second v-shaped transformation $T_{2}$ such that $h(\sigma(x),y)=g(\sigma(x),T_{2}(y))$. This conclusion extends to the more general models of first-order GARCH type with asymmetric innovation distributions.
Despite the negative conclusion, copula densities of the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ are still promising candidates for approximating or mimicking the implicit copulas of GARCH(1,1)-type models and this is the subject of the next section.
5 GARCH-mimicking copulas
5.1 Mimicking jointly symmetric GARCH copulas
There are a number of ways of finding jointly symmetric copulas that might be capable of mimicking the jointly symmetric copula of a symmetric GARCH process. We could consider:

### v2

@@ 3.2 V-transformation and inverse v-transformation of copulas
The following result, which is an adaptation of a result in McNeil, (2021), is key to constructing new copulas using v-transforms and their inverses. Part (i) is related to the result of Loaiza-Maya et al., (2018) but is expressed in a manner that involves only copulas and is free of marginal distributions.
@@
Let $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ be v-transforms with associated partial inverses $\mathcal{V}_{1}^{-1}$ and $\mathcal{V}_{2}^{-1}$, $\Delta$-functions $\Delta_{1}$ and $\Delta_{2}$ and stochastic inverses $\mathcal{V}_{1}^{\leftarrow}$ and $\mathcal{V}_{2}^{\leftarrow}$. Let $(U,V)^{\top}$ be distributed according to a copula $C$ with density $c$ and let $W_{1}$ and $W_{2}$ be iid $\mathcal{U}(0,1)$ variables, independent of $(U,V)^{\top}$.
@@
    $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$ is distributed according to a copula $C^{*}$ with density
    $c^{*}(u,v)=c(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v))\;\Delta_{1}(u)\,\Delta_{2}(v)\;+c(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v))\;\big(1-\Delta_{1}(u)\big)\,\Delta_{2}(v)\;+$
    $c(\mathcal{V}_{1}^{-1}(u),\mathcal{V}_{2}^{-1}(v)+v)\;\Delta_{1}(u)\,\big(1-\Delta_{2}(v)\big)+c(\mathcal{V}_{1}^{-1}(u)+u,\mathcal{V}_{2}^{-1}(v)+v)\;\big(1-\Delta_{1}(u)\big)\,\big(1-\Delta_{2}(v)\big).$ (13)
@@
    $(\mathcal{V}_{1}^{\leftarrow}(U,W_{1}),\mathcal{V}_{2}^{\leftarrow}(V,W_{2}))^{\top}$ is distributed according to a copula $C^{\dagger}$ with density
    $c^{\dagger}(u,v)=c(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v)).$ (14)
@@
Part (i) refers to what we call (componentwise) v-transformation of $(U,V)^{\top}$, or its distribution $C$, while part (ii) refers to what we call (componentwise independent) inverse v-transformation of $(U,V)^{\top}$, or $C$. Inverse v-transformation of $(U,V)^{\top}$ followed by v-transformation returns the original random vector $(U,V)^{\top}$ (since $\mathcal{V}(\mathcal{V}^{\leftarrow}(y,W))=y$ for $y\in[0,1]$ and any v-transform). However v-transformation of $(U,V)^{\top}$ followed by inverse v-transformation returns a different random vector with, in general, a different distribution. An exception is the following special case.
@@
Under the conditions of Theorem 1, if $C$ is jointly symmetric and $\mathcal{V}_{1}(u)=\mathcal{V}_{2}(u)=|2u-1|$ for $u\in[0,1]$, then $(\mathcal{V}_{1}^{\leftarrow}(\mathcal{V}_{1}(U),W_{1}),\mathcal{V}_{2}^{\leftarrow}(\mathcal{V}_{2}(V),W_{2}))^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,V)^{\top}$.
@@
By part (ii) of Corollary 1 the density $c$ of the copula $C$ of $(U,V)^{\top}$ can be written as $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ where $c^{*}$ is the density of the copula of $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$. By part (ii) of Theorem 1 $c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ is also the density of the inverse v-transformation of $(\mathcal{V}_{1}(U),\mathcal{V}_{2}(V))^{\top}$ with the stochastic inverses $\mathcal{V}_{1}^{\leftarrow}$ and $\mathcal{V}_{2}^{\leftarrow}$.
@@ 3.3 Implicit v-transforms in general bivariate densities
In this section we use part (i) of Theorem 1 to find a condition that allows us to write the copula density $c$ of a random vector $(Y,Z)^{\top}$ on $\mathbb{R}^{2}$ in the form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ for v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ and a copula density $c^{*}$. This result requires the concept of a v-shaped transformation on $\mathbb{R}$, which generalizes the idea found in Loaiza-Maya et al., (2018) to also allow non-symmetric transformations.
@@
A function $T$ is a v-shaped function on $\mathbb{R}$ if it can be written in the form $T(x)=T_{L}(\mu_{T}-x)$ for $x\leqslant\mu_{T}$ and $T(X)=T_{R}(x-\mu_{T})$ for $x>\mu_{T}$ where $T_{L}$ and $T_{R}$ are strictly increasing, continuous and differentiable functions on $\mathbb{R}^{+}=[0,\infty)$ such that $T_{L}(0)=T_{R}(0)$ and $\mu_{T}\in\mathbb{R}$.
@@
$\frac{f_{Y,Z}(y,z)}{f_{Y}(y)f_{Z}(z)}=g\left(T_{1}(y),T_{2}(z)\right)$ (15)
for two v-shaped transformations $T_{1}$ and $T_{2}$ on $\mathbb{R}$ and some function $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}^{+}$, then we may write $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ on $0,1]^{2}$ where $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are v-transforms given by $\mathcal{V}_{1}(u)=F_{T_{1}(Y)}\circ T_{1}\circ F_{Y}^{-1}(u)$ and $\mathcal{V}_{2}(v)=F_{T_{2}(Z)}\circ T_{2}\circ F_{Z}^{-1}(v)$ and where $c^{*}$ is the density of the copula of $(T_{1}(Y),T_{2}(Z))^{\top}$.
@@
Symmetric models of GARCH(1,1)-type provide examples of models where the density ratio for $f_{X_{1},X_{2}}$ (and also for $f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}$) may be factorized in the form (15) by taking $T_{1}(x)=T_{2}(X)=|x|$. Unfortunately, when innovations distributions are asymmetric, or when leverage is introduced into a model that is not of ARCH(1) type effect, perfect factorization of the form (15) does not seem to be possible. However, it is possible in the following case.
@@
and conclude that the copula density satisfies $c_{1}(u,v)=c_{1}^{*}(\mathcal{V}_{1}(u),|2v-1|)$ where $c_{1}^{*}$ is the copula density of $(\sigma(X_{1}),|X_{2}|)^{\top}$ or, in other words, $(\sigma_{t},|X_{t}|)^{\top}$ for any $t$. To derive $\mathcal{V}_{1}$ we must find the composition of functions $\mathcal{V}_{1}(u)=F_{\sigma}\circ\sigma\circ F_{X}^{-1}(u)$. By using the fact that the distribution function of $\sigma_{t}$ is $F_{\sigma}(s)=F_{X}\left(\sqrt{(s^{2}-\alpha_{0})/\alpha_{1}}\right)-F_{X}\left(-\sqrt{(s^{2}-\alpha_{0})(\alpha_{1}+\gamma_{1})}\right)$ we obtain after some calculation that
@@
3.4 Approaches to approximating GARCH copulas
Despite the limited number of GARCH(1,1)-type models for which we can prove that pair copulas must take the exact form $c(u,v)=c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$, this construction remains of interest as a flexible means of approximating the cross-like behaviour of the implied GARCH pair copulas. Our approach to this problem is to use simulation studies to first identify copula densities of the form $c^{*}(|2u-1|,|2v-1|)$ to approximate the jointly symmetric copulas of symmetric GARCH processes and then to consider how asymmetric v-transforms $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ can accommodate the asymmetries introduced by leverage and asymmetric innovations.
There are two ways of identifying appropriate copula densities $c^{*}$ for building models with densities of the form $c^{*}(|2u-1|,|2v-1|)$: one is the explicit method using part (ii) of Theorem 1 where we simply choose a candidate for the copula $C^{*}$ based on the insights we have collected so far; the other is the implicit method where we consider constructions that are known to yield a jointly symmetric copula and then infer the form of $c^{*}$.
Explicit choices of $C^{*}$ might be the Gumbel, Joe or 180-degree-rotated (survival) Clayton copulas. These all have upper tail dependence and may be capable of approximating the copulas of the absolute value processes of symmetric GARCH(1,1)-type models. Recall that the densities of such copulas can be thought of as blown-up versions of the upper-right quadrants of the copula densities in Figure 1. We impose one technical condition on copulas $C^{*}$ which facilitates maximum likelihood inference for models with densities of the form $c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$: we require that $c^{*}(0,0)<\infty$ which ensures that the log-likelihood is properly defined for observations at the fulcrum values.
As examples of the implicit construction we could consider:
  1. (1)
    the copula of a spherically symmetric distribution;
  2. (2)
  3. (3)
    an equal (25/25/25/25) mixture of any copula and its rotations through 90, 180 and 270 degrees.
An example of (1) is the copula of a spherical t distribution, i.e. a bivariate t distribution located at the origin with identical marginal distributions and correlation equal to zero. We call this copula the spherical t copula and denote it $C^{t}_{\nu}$. Since a spherical t-distributed random vector $(Y,Z)^{\top}$ is jointly symmetric, so is its copula. As the degree-of-freedom parameter tends to zero, this copula becomes increasingly cross-shaped and converges to an equal mixture of the comonotonicity and countermonotonicity copulas (McNeil et al.,, 2022). We denote the copula of $(|Y|,|Z|)^{\top}$ by $C^{|t|}_{\nu}$ and call it the absolute spherical t copula (ast). By (4) its density must satisfy
This one-parameter copula, which appears to be new, interpolates between independence (as $\nu\to\infty$) and comonotonicity (as $\nu\to 0$). It has upper tail dependence like the Gumbel, Joe and survival Clayton copulas. More information is collected in Section S4 in the Supplementary Material.
An example of (2) is an equal mixture of the t copulas $C^{t}_{\nu,\rho}$ and $C^{t}_{\nu,-\rho}$ for some $\rho\geqslant 0$, which is the jointly symmetric special case of a copula considered by Loaiza-Maya et al., (2018); the t mixture reduces to the spherical t copula $C^{t}_{\nu}$ when $\rho=0$. Construction (3), which is the bivariate version of a higher-dimensional family used in Oh and Patton, (2016), could be based on the Gumbel, Clayton or Joe copulas, to take advantage of the strong tail dependence of these copulas in either the upper-right or lower-left corners. The Gumbel case of (3) is the jointly symmetric special case of the convex Gumbel copula proposed by Junker and May, (2005) and used by Loaiza-Maya et al., (2018) in their analysis.
So far none of these constructions address the issue of non-exchangeability which was clearly evident in certain examples in Figure 1. A possible construction for non-exchangeable copulas was proposed by Khoudraji, (1995) and we use a version of this construction in which we generalize an exchangeable copula $C_{\theta}$ with parameter $\theta$ by setting
$C_{\theta,a_{1},a_{2}}(u,v)=C_{\theta}(u^{1-a_{1}},v^{1-a_{2}})u^{a_{1}}v^{a_{2}}$ (17)
for $a_{1}\in[0,1]$ and $a_{2}\in[0,1]$. The resulting copula is non-exchangeable when $a_{1}\neq a_{2}$. We apply the Khoudraji extension to copulas before either mixing them with their rotations or applying inverse v-transformation.
3.5 Simulation study of the approximation of $C_{1}$
To get some insight into the approximations that work best for the pair copulas implied by GARCH-type models, we carry out a simulation study in which we focus on estimation of the first-order copula $C_{1}$. Samples of 50000 values are generated from ARCH(1)-type and GARCH(1,1)-type processes using the parameter values and innovation distributions summarized in Table 1. The copula $C_{1}$ is estimated by fitting a first-order D-vine to the data using the two-stage method developed by Chen and Fan, (2006) in which the marginal distribution of the process is estimated by the rescaled empirical distribution function (EDF) to produce pseudo-copula data. The explicit and implicit constructions of Section 3.4 are tested and fitted models are compared according to their AIC values. Note that, the first-order copulas $C_{1}$ are the only pair copulas for which we can produce pseudo-copula data using only the rescaled EDF; we address approximations for higher-order copulas $(C_{k})_{k>1}$ indirectly in Section 4.2 by fitting higher-order simplified D-vines to simulated GARCH data.
Under the assumption that the model is first-order Markov with known copula, Chen and Fan, (2006) show consistency and asymptotic normality of the resulting parameter estimates subject to technical conditions which include that the process should be $\beta$-mixing at a geometric rate. All the GARCH(1,1)-type processes we consider in our simulation satisfy this mixing condition (Francq and Zakoïan,, 2006). However, the first-order Markov assumption is only true for ARCH(1)-type processes and our models contain approximating copulas rather than correctly specified copulas. Nagler et al., (2022) supply some very general results which can be applied to semiparametric inference for simplified D-vines of higher order when the pair copulas are misspecified. Their results suggest that consistency and asymptotic normality of the procedure hold in the misspecified model, albeit convergence is to the parameter values that solve the model score equations, which coincide with the true values in a correctly specified model.
3.5.1 Jointly symmetric models
Model ARCH(1) GARCH(1,1)
Copula (form of density) $p$ | $(\epsilon_{t})$ Gauss t4 t2.5 Gauss t4 t2.5
$c^{t}_{\nu}(u,v)=c^{|t|}_{\nu}(|2u-1|,|2v-1|)$ 1 528 0 126 103 7 95
$\tfrac{1}{2}\left(c^{t}_{\rho,\nu}(u,v)+c^{t}_{-\rho,\nu}(u,v)\right)$ 2 530 2 128 101 9 97
$\tfrac{1}{4}\left(c_{\theta}^{\text{Cl}}(u,v)+c_{\theta}^{\text{Cl},90}(u,v)+c_{\theta}^{\text{Cl},180}(u,v)+c_{\theta}^{\text{Cl},270}(u,v)\right)$ 1 1676 426 189 491 291 145
$\tfrac{1}{4}\left(c_{\theta}^{\text{Gu}}(u,v)+c_{\theta}^{\text{Gu},90}(u,v)+c_{\theta}^{\text{Gu},180}(u,v)+c_{\theta}^{\text{Gu},270}(u,v)\right)$ 1 913 150 182 228 120 156
$\tfrac{1}{4}\left(c_{\theta}^{\text{J}}(u,v)+c_{\theta}^{\text{J},90}(u,v)+c_{\theta}^{\text{J},180}(u,v)+c_{\theta}^{\text{J},270}(u,v)\right)$ 1 1982 593 276 666 431 232
$c_{\theta}^{\text{Cl},180}(|2u-1|,|2v-1|)$ 1 529 146 292 174 105 216
$c_{\theta}^{\text{Gu}}(|2u-1|,|2v-1|)$ 1 1167 444 379 422 307 290
$c_{\theta}^{\text{J}}(|2u-1|,|2v-1|)$ 1 650 50 145 144 65 131
$\tfrac{1}{4}\left(c_{\theta,a_{1},a_{2}}^{\text{Cl}}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Cl},90}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Cl},180}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Cl},270}(u,v)\right)$ 3 1170 430 102 385 288 70
$\tfrac{1}{4}\left(c_{\theta,a_{1},a_{2}}^{\text{Gu}}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Gu},90}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Gu},180}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{Gu},270}(u,v)\right)$ 3 854 154 66 199 124 64
$\tfrac{1}{4}\left(c_{\theta,a_{1},a_{2}}^{\text{J}}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{J},90}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{J},180}(u,v)+c_{\theta,a_{1},a_{2}}^{\text{J},270}(u,v)\right)$ 3 1680 597 178 632 435 151
$c_{\theta,a_{1},a_{2}}^{\text{Cl},180}(|2u-1|,|2v-1|)$ 3 46 147 156 89 98 119
$c_{\theta,a_{1},a_{2}}^{\text{Gu}}(|2u-1|,|2v-1|)$ 3 598 448 221 330 303 159
$c_{\theta,a_{1},a_{2}}^{\text{J}}(|2u-1|,|2v-1|)$ 3 93 54 11 47 63 18
$c_{\nu,a_{1},a_{2}}^{|t|}(|2u-1|,|2v-1|)$ 3 0 3 0 0 0 0
Table 2: Comparison of approximating copulas for the first-order copula of symmetric ARCH(1) and GARCH(1,1) processes. Estimates are obtained by fitting first-order D-vines using semiparametric method of Chen and Fan, (2006). Each column shows the differences between AIC and the minimum value of AIC for that column; thus a 0 indicates the model with lowest AIC. The superscripts Cl, Gu and J refer to Clayton, Gumbel and Joe copulas while 90, 180 and 270 denote rotations. Densities $c_{\theta,a_{1},a_{2}}$ correspond to the Khoudraji construction in (17).
The results in Table 2 relate to classical ARCH(1) and GARCH(1,1) models with symmetric innovations. The innovation distributions considered are Gaussian, Student t4 and Student t2.5, giving six different datasets in total. The spherical t copula $C^{t}_{\nu}$ works quite well in all cases (despite its exchangeability) and is the favoured model for the case of ARCH(1) with t4 innovations. It is notable that the spherical t copula generally outperforms the mixture of two t copulas; in the latter the estimate of $\rho$ is very close to zero suggesting the model is poorly specified for this application. For the other datasets the favoured copula has density of the form $c^{*}(|2u-1|,|2v-1|)$ where $C^{*}$ is the Khoudraji extension (17) of the ast copula; the second-best copula is the model where $C^{*}$ is the Khoudraji extension of either the Joe copula (four cases) or the survival Clayton copula (one case). Models constructed using explicit inverse v-transformation are more effective than mixture copulas constructed according to approaches (2) and (3) of Section 3.4. While the Gumbel copula gives the best mixtures, the Joe and survival Clayton copulas give better inverse-v-transformed models than the Gumbel copula.
3.5.2 Models with asymmetries
When models contain asymmetric innovations, and potentially also leverage, we consider approximating GARCH pair copulas using the construction $c^{*}(\mathcal{V}_{1}(u),\mathcal{V}_{2}(v))$ with asymmetric v-transforms $\mathcal{V}(u;\delta,\kappa)$ taken from the family (12), which includes the linear special case $\mathcal{V}(u;\delta)$. Section S5.2 of the Supplementary Material contains extensive results for ARCH(1) and GARCH(1,1) models with skewed t innovations, with and without leverage effects. The most important observation in these analyses is that, when two different asymmetric v-transforms $\mathcal{V}_{1}(u)$ and $\mathcal{V}_{2}(v)$ are used, the Khoudraji extension of the copula $C^{*}$ adds very little. This is a helpful insight from a computational point of view because evaluating the density of the Khoudraji extension requires, not only the density $c^{*}$ of the parent copula, but also the function $C^{*}$ itself, leading to much slower evaluation (particularly in the ast case). The best models are of the two-parameter form $c^{*}(\mathcal{V}(u;\delta_{1},\kappa_{1}),\mathcal{V}(v;\delta_{2},\kappa_{2}))$ with $c^{*}$ taken to be $c^{|t|}_{\nu}$ or $c^{\text{J}}_{\theta}$. However models of the form $c^{*}(\mathcal{V}(u;\delta_{1}),\mathcal{V}(v;\delta_{2}))$ with two asymmetric linear v-transforms already offer a large improvement on jointly symmetric models and offer the advantages of greater parsimony and greater tractability, as we discuss in Section 4.1. Moreover, the analyses of Section 4.2 will show that, in the presence of asymmetry, models with two asymmetric linear v-transforms tend to give better estimates of $C_{1}$ than mixtures of t copulas or convex Gumbel copulas.
4 Simplified D-vines for stochastic volatility
4.1 Proposal for a simplified D-vine structure
In this section we use insights from Sections 2 and 3 to propose a class of simplified D-vines that behave like GARCH(1,1) processes and offer potential alternative models for financial time series. A guiding principle in constructing our models is parsimony. GARCH(1,1) models have relatively few parameters and an effective model that incorporates both leverage and skewed innovations can be created with six parameters ($\alpha_{0}$, $\alpha_{1}$, $\beta_{1}$, $\gamma_{1}$, $\xi$, $\lambda$). The first four control the volatility equation and the final two describe the shape and skew of the innovation distribution; a seventh parameter $\mu$ could be added for a non-zero mean. With a handful of parameters, the GARCH(1,1) model is able to model the persistence of volatility observed in financial data in a way that low-order ARCH($p$) models are not. To achieve a similar effect in D-vine models, we construct sequences of pair copulas $(C_{k})_{k\in\mathbb{N}}$ using three main ideas.